# Distraction is All You Need for Multimodal Large Language Model Jailbreaking

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) bridge the gap between visual and textual data, enabling a range of advanced applications. However, complex internal interactions among visual elements and their alignment with text can introduce vulnerabilities, which may be exploited to bypass safety mechanisms. To address this, we analyze the relationship between image content and task and find that the complexity of subimages, rather than their content, is key. Building on this insight, we propose the Distraction Hypothesis, followed by a novel framework called Contrasting Subimage Distraction Jailbreaking (CS-DJ), to achieve jailbreaking by disrupting MLLMs alignment through multi-level distraction strategies. CS-DJ consists of two components: structured distraction, achieved through query decomposition that induces a distributional shift by fragmenting harmful prompts into sub-queries, and visual-enhanced distraction, realized by constructing contrasting subimages to disrupt the interactions among visual elements within the model. This dual strategy disperses the model’s attention, reducing its ability to detect and mitigate harmful content. Extensive experiments across five representative scenarios and four popular closed-source MLLMs, including GPT-4o-mini, GPT-4o, GPT-4V, and Gemini-1.5-Flash, demonstrate that CS-DJ achieves average success rates of 52.40% for the attack success rate and 74.10% for the ensemble attack success rate. These results reveal the potential of distraction-based approaches to exploit and bypass MLLMs’ defenses, offering new insights for attack strategies.
@@ 2.1 Multimodal Large Language Models
In recent years, Multimodal Large Language Models (MLLMs) [34] have emerged as the leading solution for handling complex multimodal data, such as images[43] and audio [5]. Inspired by instruction-tuning algorithms in LLMs, many MLLMs are fine-tuned from pre-trained LLMs. Notable examples include LLaVA [13] and MiniGPT-4 [46], which utilize similar architectures by connecting pre-trained vision backbones to large language models via a linear layer. These models are then fine-tuned using ChatGPT-generated data, enabling deep integration between vision and language. To enhance performance and usability, LLaVA-1.5 [12] replaces the linear layer with MLPs and incorporates prompt-based response formats optimized for academic VQA datasets.
MM1 [20] and MoE-LLaVA [11] further introduce MoE [26] structure to achieve better performance, without increasing computational cost. However, the integration of emerging modalities further intensifies the challenges associated with the safety alignment.
To achieve this, MLLMs are typically fine-tuned using data annotated with human preferences to ensure that their outputs are helpful and harmless. Despite these efforts, this approach remains vulnerable to attacks using OOD inputs, which will be thoroughly explored in this work.
2.2 Jailbreak Attacks against MLLMs
Jailbreak attacks against MLLMs can be conducted through both visual and textual inputs, leading to complex and diverse attack patterns [8, 21]. Current jailbreak attack techniques can be broadly categorized into two types: image perturbation injection and prompt-to-image infection jailbreaks. Among them, image perturbation injection techniques are particularly effective for white-box jailbreak scenarios.
In particular, to bypass the model’s security detection mechanisms, Qi et al. [23] conceal harmful textual information within images by injecting invisible noise [45] through gradient optimization. Furthermore, JIP [25] conceals harmful image information within benign images through latent space semantic alignment. To enable jailbreaks on black-box MLLMs, Hades [10] investigates the influence of image harmfulness on model jailbreaks and proposes a prompt-to-image infection jailbreak attack method. It generates harmful images using diffusion models [27, 39, 19] and amplifies their effect through optimization. The generated image is then combined with the query text to form a jailbreak prompt [15, 17, 6], successfully achieving a jailbreak. In a word, existing MLLM jailbreak attack methods generally focus on crafting out-of-distribution (OOD) inputs. However, it remains unclear how to effectively leverage the visual space to construct OOD inputs for successful black-box MLLM jailbreaks. To address this, we propose CS-DJ, which employs distraction to guide the construction of OOD inputs.
@@ 4.1 Experiment Setup
Dataset. The datasets are primarily utilized to provide or construct visual inputs tailored to the requirements of different methods. For evaluating Hades, we followed the experimental setting outlined in its original paper, utilizing the Hades dataset [10] to supply the visual inputs. This dataset comprises five representative categories related to real-world visual information: (1) Violence, (2) Financial, (3) Privacy, (4) Self-Harm, and (5) Animal. Each category contains 150 queries, resulting in a total of 750 harmful queries. For evaluating CS-DJ, we employed LLaVA-CC3M-Pretrain595K  [^1] https://huggingface.co/datasets/liuhaotian/LLaVA-CC3M-Pretrain-595K to retrieve the contrasting images. To improve retrieval efficiency, we randomly selected a subset of 10,000 images for the retrieval dataset, using a different random seed for each experiment to ensure diversity. All harmful queries used throughout the experiments were derived from the Hades dataset.
Victim Models. In our main experiments, We evaluated the performance of the four most popular closed-source MLLMs: GPT-4o-Mini, GPT-4o, GPT-4V [1], and Gemini-1.5-Flash [30]. The specific versions are as follows: GPT-4o: GPT-4o-2024-08-06, GPT-4o-Mini: GPT-4o-Mini-2024-07-18, GPT-4V: GPT-4-1106-vision-preview, and Gemini-1.5-flash: Gemini-1.5-flash-001  [^2] Since Gemini-1.0-Pro-Vision has been deprecated on July 12, 2024, we used its upgraded version, Gemini-1.5-Flash..
Evaluation Metrics. To assess the effectiveness of our proposed CS-DJ framework, we employ two primary evaluation metrics: Attack Success Rate (ASR) and Ensemble Attack Success Rate (EASR). ASR quantifies the proportion of successful jailbreak attempts by determining whether the model’s response meets predefined conditions for harmful content. This determination is guided by a security discriminator that classifies responses as harmful or benign based on specific criteria. We utilize Beaver-Dam-7b [7], trained on high-quality human feedback data, as the security discriminator, leveraging its ability to detect both the harmfulness and helpfulness of responses to their corresponding queries. This ensures a robust and consistent evaluation of attack success across various scenarios. EASR [40] evaluates the effectiveness of a set of jailbreak templates by measuring the proportion of queries for which at least one template within the subset successfully bypasses the target MLLM’s defenses. While ASR focuses on the success rate of an individual template, EASR provides a broader assessment by considering the combined success of a group of templates, offering deeper insights into the MLLM’s susceptibility to jailbreak attempts.
@@ 4.2 Main Experiment
A detailed analysis of Hades’ results across the four models reveals that GPT-4V outperforms the other three models in both ASR and EASR. In contrast, the results of CS-DJ indicate that Gemini-1.5-Flash achieved the highest performance. These findings suggest that the impact of visual input harmfulness and distraction varies across different models. In a word, the results demonstrate that distracting the model’s attention is a more effective strategy for enhancing jailbreak success rates.
@@ 4.4 Impact of Contrasting Visuals
To validate the effectiveness of contrasting visuals, we conduct experiments from two perspectives: (1) the impact of distraction induced by varying the number of subimages, and (2) the effect of inter-subimage distraction with a fixed number of subimages. Since altering the number of subimages directly affects the layout of the final composite image, we maintain a fixed column count of 3 and adjust the number of rows to control the number of subimages. This strategy reduces distortion due to excessive changes in aspect ratio and mitigates potential limitations in the MLLM’s image encoding capabilities. We conduct different experiments with various configurations, involving 0, 3, 6, 9, and 12 Contrasting visual SubImages, corresponding to 3SQ, 3SQ+3CSI, 3SQ+6CSI, 3SQ+9CSI, and 3SQ+12CSI, respectively. Furthermore, to explore the influence of inter-subimage distraction on jailbreak success rates, we conduct experiments using three settings: retrieving 9 most contrasting images, retrieving 9 most similar images, and using a single, most similar image from the dataset for all 9 subimages, denoting 3SQ+9CSI, 3SQ+9SSI, and 3SQ+9SinSI, respectively.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) bridge the gap between visual and textual data, enabling a range of advanced applications. However, complex internal interactions among visual elements and their alignment with text can introduce vulnerabilities, which may be exploited to bypass safety mechanisms. To address this, we analyze the relationship between image content and task and find that the complexity of subimages, rather than their content, is key. Building on this insight, we propose the Distraction Hypothesis, followed by a novel framework called Contrasting Subimage Distraction Jailbreaking (CS-DJ), to achieve jailbreaking by disrupting MLLMs alignment through multi-level distraction strategies. CS-DJ consists of two components: structured distraction, achieved through query decomposition that induces a distributional shift by fragmenting harmful prompts into sub-queries, and visual-enhanced distraction, realized by constructing contrasting subimages to disrupt the interactions among visual elements within the model. This dual strategy disperses the model’s attention, reducing its ability to detect and mitigate harmful content. Extensive experiments across five representative scenarios and four popular closed-source MLLMs, including GPT-4o-mini, GPT-4o, GPT-4V, and Gemini-1.5-Flash, demonstrate that CS-DJ achieves average success rates of 52.40% for the attack success rate and 74.10% for the ensemble attack success rate. These results reveal the potential of distraction-based approaches to exploit and bypass MLLMs’ defenses, offering new insights for attack strategies. Our code is available at https://github.com/TeamPigeonLab/CS-DJ.
@@ 2.1 Multimodal Large Language Models
Recently, Multimodal Large Language Models (MLLMs) [31] have become a prominent approach for processing and understanding heterogeneous multimodal data, including visual[39] and auditory [5] modalities. Building on instruction-tuning techniques developed for LLMs, most MLLMs are adapted from pre-trained LLMs through fine-tuning. Notable examples include LLaVA [14] and MiniGPT-4 [41], which share a similar architectural design by integrating pre-trained vision backbones with large language models via a projection layer, enabling effective vision-language alignment. Building upon this foundation, LLaVA-1.5 [13] further enhances its architecture by adopting a two-layer MLP for better vision-language mapping and increasing the vision encoder’s input resolution for improved visual understanding. However, the incorporation of emerging modalities further intensifies the challenges of ensuring robust safety alignment in MLLMs. To address this challenge, advanced MLLMs, such as GPT-4o [7] and Gemini 1.5-Flash [27], employ a combination of alignment strategies, including post-training methods and integrated classifiers, to enhance both helpfulness and safety. Despite these efforts, MLLMs remains vulnerable to attacks using OOD inputs, which will be explored in this work.
2.2 Jailbreak Attacks on MLLMs
Jailbreak attacks against MLLM can be performed through visual and textual input, leading to complex and diverse attack patterns [9, 20]. As a result, numerous jailbreak methods [22, 6, 17, 12, 24] have been continuously proposed. Qi et al. [22] used visual adversarial examples to bypass the security mechanisms of visual-textual MLLMs. However, this type of attack [20, 22] is typically only effective in white-box environments and is often difficult to apply to state-of-the-art black-box models. FigStep [6] introduced a simple yet effective black-box jailbreak approach that circumvents safety alignment by trasfering prohibited content as images using typographic transformations. MM-SafetyBench [17] utilized a multi-step approach to construct jailbreak input, incorporating question generation, rephrasing, extraction of unsafe key phrases, and query-to-image conversion. Furthermore, Hades [12] investigated the influence of image harmfulness on model jailbreaks and proposed a prompt-embedded image infection jailbreak attack method. It generated harmful images using diffusion models and amplified the jailbreak effect through optimization. Unlike these works, CS-DJ explores the safety of MLLMs from a new perspective—visual distraction.
@@ 4.1 Experiment Setup
Dataset. The datasets are primarily utilized to provide or construct visual inputs tailored to the requirements of different methods. For evaluating Hades, we followed the experimental setting outlined in its original paper, utilizing the Hades dataset [12] to supply the visual inputs. This dataset comprises five representative categories related to real-world visual information: (1) Violence, (2) Financial, (3) Privacy, (4) Self-Harm, and (5) Animal. Each category contains 150 queries, resulting in a total of 750 harmful queries. For evaluating CS-DJ, we employed LLaVA-CC3M-Pretrain595K https://huggingface.co/datasets/liuhaotian/LLaVA-CC3M-Pretrain-595K to retrieve the contrasting images. To improve retrieval efficiency, we randomly selected a subset of 10,000 images for the retrieval dataset, using a different random seed for each experiment to ensure diversity. All harmful queries used throughout the experiments were derived from the Hades dataset.
Victim Models. In our main experiments, We evaluated the performance of the four most popular closed-source MLLMs: GPT-4o-Mini, GPT-4o, GPT-4V [1], and Gemini-1.5-Flash [27]. The specific versions are as follows: GPT-4o: GPT-4o-2024-08-06, GPT-4o-Mini: GPT-4o-Mini-2024-07-18, GPT-4V: GPT-4-1106-vision-preview, and Gemini-1.5-flash: Gemini-1.5-flash-001 Since Gemini-1.0-Pro-Vision has been deprecated on July 12, 2024, we used its upgraded version, Gemini-1.5-Flash..
Evaluation Metrics. To assess CS-DJ, we employ two primary evaluation metrics: Attack Success Rate (ASR) and Ensemble Attack Success Rate (EASR). ASR [6, 12] quantifies the proportion of successful jailbreak attempts by assessing whether a model’s response meets predefined harmfulness criteria. The harmfulness criteria come from the security discriminator Beaver-Dam-7b [8], used to detect both the harmfulness and helpfulness of responses to their corresponding queries. EASR [36] measures the effectiveness of jailbreak templates by calculating the proportion of queries where at least one template successfully bypasses the target MLLM’s defenses. While ASR focuses on the success rate of individual templates, EASR measures the combined success rate of a group of templates.
@@ 4.2 Main Experiment
A detailed analysis of Hades’ results across the four models reveals that GPT-4V outperforms the other three models in both ASR and EASR. In contrast, the results of CS-DJ indicate that Gemini-1.5-Flash achieved the highest performance. These findings suggest that the impact of visual input harmfulness and distraction varies across different models. In a word, the results demonstrate that distracting the model’s attention is a more effective strategy for enhancing jailbreak success rates. More quantitative and qualitative experimental results, including those tested on open-source models, are provided in the appendix.
@@ 4.4 Impact of Contrasting Visuals
To validate the effectiveness of contrasting visuals, we conduct experiments from two perspectives: (1) the impact of distraction induced by varying the number of subimages, and (2) the effect of inter-subimage distraction with a fixed number of subimages. Since altering the number of subimages directly affects the layout of the final composite image, we maintain a fixed column count of 3 and adjust the number of rows to control the number of subimages. This strategy reduces distortion due to excessive changes in aspect ratio and mitigates potential limitations in the MLLM’s image encoding capabilities. We conduct different experiments with various configurations, involving 0, 3, 6, 9, and 12 Contrasting visual SubImages, corresponding to 3SQ, 3SQ+3CSI, 3SQ+6CSI, 3SQ+9CSI, and 3SQ+12CSI, respectively. To explore the influence of inter-subimage distraction on jailbreak success rates, we conduct experiments using three settings: retrieving 9 most contrasting images, retrieving 9 most similar images, and using a single, most similar image from the dataset for all 9 subimages, denoting 3SQ+9CSI, 3SQ+9SSI, and 3SQ+9SinSI, respectively.
@@ 5 Conclusion
Acknowledgment. This research is supported by Guangzhou Basic and Applied Basic Research Project (No: 2025A04J3536, 2025A03J3128, 2024A03J0324), Guangzhou University Research Project (No: RQ2021013), Guangdong Basic and Applied Basic Research Foundation (No: 2023A1515110077), NSFC (No: 62402131), and the Postdoctoral Fellowship Program of CPSF (No: GZC20230595).
Supplementary Material
In this section, we first provide the theoretical foundation of distraction hypothesis in Section 6 and additional details of the experimental settings in Section 7. Then, in Section 8, we present the prompts used by CS-DJ during the query decomposition and jailbreaking execution phases. Finally, Section 9 shows more quantitative and qualitative experimental results, including those tested on open-source model.
6 Theoretical Foundation of Distraction Hypothesis
The Distraction Hypothesis is grounded in the theoretical idea that semantically diverse and locally inconsistent subimages disrupt the model’s attention and semantic coherence, leading to Semantic OOD (SOOD). These inputs deviate from the MLLM’s learned distribution, causing it to struggle with processing them, which ultimately leads to a degradation of its defenses. In detail, since RLHF-trained MLLMs on safety-aligned datasets are expected to generate only appropriate responses, we follow existing OOD detection frameworks in image classification and define an input as SOOD if it leads to an undesired output, such as harmful content, off-topic responses, or unjustified refusals. Formally, if the model’s output $f(x_{\text{new}})$ falls into the set of undesired outputs $C_{\text{SOOD}}$, the input $x_{\text{new}}$ is classified as SOOD: $f(x_{\text{new}})\in C_{\text{SOOD}}$. This reflects the model’s inability to handle such inputs, marking them as SOOD. Specifically, this work focuses on SOOD leading to harmful outputs. Unlike previous adversarial attacks, which rely on injecting low-level visual noise and require extensive optimization with limited transferability, our approach leverages semantic-level distraction to exploit model vulnerabilities. Our findings lay a foundation for future research on understanding and mitigating SOOD-based threats. Given the black-box setting of advanced MLLMs, the degree of SOOD cannot be directly measured. Therefore, we introduce Distraction Distance (Eqn. (8)) to estimate the SOOD degree. However, the theory is still in development, requiring further refinement.
7 Experimental Setting
Here, we provide a detailed overview of the experimental settings. The same configuration was applied to both Hades and CS-DJ. For the GPT series models, the temperature was set to 0.1, with a maximum output length of 1000 tokens. For the Gemini series models, the temperature was adjusted to 0.2, with the maximum output length increased to 2048 tokens. Additionally, each subimage grid was rendered at a resolution of 500 $\ast$ 500 pixels. If a retrieved image had a resolution smaller than this size, it was placed at the center of the grid without resizing. Otherwise, the image was scaled proportionally to fit within a 500-pixel width or height while preserving its aspect ratio. Sub-queries were converted into images using the Super Moods font, with the font color set to red and a font size of 50.
8 Prompt Design
8.1 Decomposition Query
CS-DJ employs an auxiliary decomposition model $\mathcal{G}$ to break down the raw query into multiple sub-queries, thereby introducing structured distraction. The prompt for decomposing into three sub-queries is provided in Figure 4. Notably, the decomposition by $\mathcal{G}$ is considered complete only when the generated responses strictly adhere to the specified format. Otherwise, the process is retried up to a maximum of 5 attempts in practice.
Figure 4: Prompt for query decomposition. specifically, the placeholder {jailbreak query} is replaced with the raw query. The bolded portion should be modified to align with the number of sub-queries.
8.2 Jailbreaking Execution
Figure 5 illustrates the prompt used for the jailbreaking execution of CS-DJ, carefully designed to enhance the distraction effect. The instruction consists of three main sections: the role-guiding section (in red), the task-guiding section (in black), and the visual-guiding section (in blue). The role-guiding section establishes a scenario for the model, providing the contextual framework for the subsequent tasks. The task-guiding section instructs the model to simultaneously perform multiple tasks within specific subimages, increasing complexity and deliberately dispersing its focus across different objectives. Lastly, the visual-guiding section introduces misleading cues, implying that other subimages might be useful, further diverting the model’s attention.
Figure 5: Prompt for the jailbreaking execution. The bolded portion should be modified to align with the number of subimages.
9 Additional Experimental Results
9.1 Open-source Model Results
To further validate CS-DJ, we conducted experiments based on the open-source MLLM, LLaVA-OneVision-Chat-7B [11], in a single round. Results in Table 6 show a $7.07\%$ increase in ASR for CS-DJ over Hades, suggesting that closed-source models have undergone more effective safety alignment. The Distraction mechanism also outperforms methods enhancing visual harmfulness in bypassing safety detection.
Method Ani. Fin. Priv. Self-H. Viol. ASR (%)
Hades 14.00 72.00 64.67 24.00 69.33 48.80
CS-DJ 19.33 83.33 65.33 40.00 71.33 55.87
Table 6: ASR result of CS-DJ and Hades on LLaVA-OneVision-Chat-7 across different categories.
9.2 Visualization Comparison
We analyzed the attention maps of both methods during the generation of the first token using LLaVA-OneVision-Chat-7B [11]. As shown in Figure 6, Hades focuses on the harmful content in the image, while CS-DJ effectively disperses the model’s attention, supporting the role of distraction in bypassing model defenses.
Figure 6: Comparative Attention Visualization of Hades and CS-DJ.
9.3 Additional Baselines
We evaluated two additional representative models, FigStep [6] and MM-SafetyBench [17], using the Beaver-Dam-7b [8] evaluation. Shown in table 7, all experiments were conducted in a single round, based on GPT-4o-2024-08-06. The ASR results demonstrate that CS-DJ outperforms the three baseline models.
Method Ani. Fin. Priv. Self-H. Viol. ASR (%)
Figstep 2.66 4.00 3.33 0 8.66 3.73
MM-SafetyBench 3.33 4.00 8.66 2.66 10.66 5.86
CS-DJ 19.33 73.33 56.66 18.00 51.33 43.73
Table 7: ASR of CS-DJ and other baselines.
9.4 Additional Jailbreak Cases
This section presents detailed jailbreak cases for GPT-4o-min, GPT-4o, GPT-4V, and Gemini-1.5-Flash, as shown in Figure 7, Figure 8, Figure 9, and Figure 10, respectively.
Figure 7: Jailbreak case of GPT-4o-mini.
Figure 8: Jailbreak case of GPT-4o.
Figure 9: Jailbreak case of GPT-4V.
Figure 10: Jailbreak case of Gemini-1.5-Flash.